OpenStudy (anonymous):

Could anyone check my working out the Fourier series for \[ f(t)=\bigl(\begin{smallmatrix} 1 ...... 0

OpenStudy (anonymous):

The bit cut off was \[[-\frac{1}{nw}\cos(nw(T/2))]-[[-\frac{1}{nw}\cos(nw(T/2))]-[-\frac{1}{nw}\cos(nwT)]]\] \[=-[\frac{1}{nw}\cos(nw(T/2))]+[\frac{1}{nw}\cos(nw(T/2))]-[\frac{1}{nw}\cos(nwT)]]\]

OpenStudy (anonymous):

\[=-[\frac{1}{nw}\cos(n\frac{2 \pi T}{2T})]+[\frac{1}{nw}\cos(n\frac{2 \pi T}{2T}))]-[\frac{1}{nw}\cos(n\frac{2 \pi T}{T})]]\]\[=-[\frac{1}{nw}\cos(n\pi)]+[\frac{1}{nw}\cos(n\pi))]-[\frac{1}{nw}\cos(n2 \pi)]]\]

OpenStudy (experimentx):

table after (19) http://mathworld.wolfram.com/FourierSeries.html

OpenStudy (experimentx):

here is better example for square wave http://mathworld.wolfram.com/FourierSeriesSquareWave.html

OpenStudy (anonymous):

Damn! I left in an unnecessary bracket!

OpenStudy (anonymous):